A 6-foot y’all tent casts a 9-foot shadow. If the tent is next to a cardboard box and the box casts a 6-foot long shadow, how tall is the box?
1
Expert's answer
2019-12-06T09:22:59-0500

Let us say the height of cardboard box is h feet


and the angle of elevation = "\theta"


"\tan\theta" = "\frac{Height\thinspace of\thinspace the\thinspace tent }{Length\thinspace of\thinspace tent's\thinspace shadow}" = "\frac{Height\thinspace of\thinspace the\thinspace box }{Length\thinspace of\thinspace box's\thinspace shadow}"


"\implies" "\frac{6 }{9}" = "\frac{h }{6}"


"9h" = "36"


So "h=4" feet


Answer: the height of the cardboard box is 4 feet.